OpenStudy (anonymous):

consider the ellipse : 9x^2 +2y^2=18 I have found everything except the endpoints of the minor axis.

OpenStudy (amistre64):

its the same precedure as the endpoints of the major axis, just with the smaller denominator

OpenStudy (amistre64):

\[ 9x^2 +2y^2=18\] \[ \frac{9x^2}{18} +\frac{2y^2}{18}=\frac{18}{18}\] \[ \frac{x^2}{18/9} +\frac{y^2}{18/2}=1\] \[ \frac{x^2}{2} +\frac{y^2}{9}=1\]

OpenStudy (amistre64):

the denoms i believe refer to "a" and "b" by convention, where a is the larger/major axis length; and b is the smaller

OpenStudy (anonymous):

thats what I got but I don't know what to do about the 2... you need to square it such as 3^2 = 9

OpenStudy (amistre64):

b^2 = 2 b = sqrt(2)

OpenStudy (amistre64):

since the smaller denom is under the x, the minor axis is parallel to the x axis

OpenStudy (amistre64):

and this thing is centered at the origin; so the minor end points are (sqrt(2),0) and (-sqrt(2),0)

OpenStudy (anonymous):

ohhhhhh ! duh!!! i completely forgot that you can do that! thanks so much it was killing me!!!

OpenStudy (amistre64):

:) youre welcome, good luck